The angle of elevation of the top of a tower 30 m high, from two points on the level ground on its opposite sides are 45 degrees and 60 degrees. What is the distance between the two points?

Grade:

1 Answers

raj kaushik
15 years ago

CHECK OUT THIS SOLUTION  !!!! 

Let OT be te tower.



Therefore, Height of tower = OT = 30 m
Let A and B be the two points on the level ground on the opposite side of tower OT.

Then,

angle of elevation from A = Trigonometric RatiosTAO = 45o
and angle of elevation from B = Trigonometric RatiosTBO = 60o
Distance between AB = AO + OB = x + y (say)

Now, in right triangle ATO,
tan 45o = OT/AO =30/x         
=> x =  30/tan
45o  = 30 m

and in right traingle BTO
tan 60o =  OT/OB =30/y         

=> y = Trigonometric identities = 17.32 m

Hence, the required distance = x + y = 30 + 17.32 = 47.32 m